OpenStudy (anonymous):

The legs of a right triangle are 3 units and 6 units. What is the length of the hypotenuse? Round your answer to the nearest hundredth. 5.20 units 6.71 units 9.00 units 18.00 units

OpenStudy (jackofalltradez):

Use the Pythagorem Theorem: a^2 + b^2 = c^2

OpenStudy (matt101):

Just use the Pythagorean Theorem: \[a^2+b^2=c^2\] Where c is the hypotenuse and a and b are the other two sides.

OpenStudy (anonymous):

so would the answer be a or b ? im stuck between the to >.<

OpenStudy (jackofalltradez):